OpenStudy (anonymous):

The length of each side of a cube is multiplied by 3. What is the change in the surface area of the cube?

OpenStudy (anonymous):

the surace area of a cube of edge x is SA = 6*x^2 now if the sides are multiplied by 3, then the new SA_new = 6*(3x)^2 = 6*9x^2 = 9*(6x^2). the new surface area is 9 times bigger than the old one.